OpenStudy (anonymous):

What is the axis of symmetry for the graph of y = 2x2 − 8x + 10?

OpenStudy (anonymous):

. x = 2 b. x = −2 c. x = 4 d. x = −4

OpenStudy (alexandervonhumboldt2):

use the formula \[x=-\frac{ b }{ 2a }\] where ax^2+bx+c=0

OpenStudy (alexandervonhumboldt2):

in your case, a=2, b=-8 and c=10

OpenStudy (anonymous):

isnt it -b/2a?

OpenStudy (alexandervonhumboldt2):

it is same

OpenStudy (alexandervonhumboldt2):

it doesnt matter where the minus is

OpenStudy (anonymous):

i got 16/4

OpenStudy (anonymous):

so x = 4

OpenStudy (alexandervonhumboldt2):

no: \[-\frac{ b }{ 2a }=\frac{ -(-8) }{ 2*2 }=\frac{ 8 }{ 4 }=??\]

OpenStudy (anonymous):

oh oops. i did 8 x 2

OpenStudy (alexandervonhumboldt2):

yeah

OpenStudy (anonymous):

x=2

OpenStudy (alexandervonhumboldt2):

correct

OpenStudy (anonymous):

TYVM!